Here are the results of a cross-tabulation table drawn from sample data. Two variables were measured in this study, age and concern about rising inequality in US society.



Very Somewhat Not at all

concerned concerned concerned

Over 30 10 35 55


30 or younger 60 25 15


n =

  1. In our study of feelings about inequality in the US ______________% of the sample is over 30 and ____________ % is 30 or younger.
  2. According to our study____________ % of respondents are very concerned about rising inequality in the US, while __________ % of people we surveyed are not at all concerned.
  3. Importantly, we note that ___________% of respondents who are 30 or younger are very concerned with rising inequality, while only __________% of respondents over 30 feel this way about the issue.
1
Expert's answer
2021-03-10T04:44:56-0500

"1)\frac{100}{200}\times 100=50, \frac{100}{200}\times 100=50\\
2)\frac{70}{200}\times 100=35, \frac{70}{200}\times 100=35\\
3)\frac{60}{100}\times 100=60, \frac{10}{100}\times 100=50\\"